Exercice 6. Let be (Xi,..., Xn) an iid sample from the Bernoulli distribution with parameter θ, ie. I. What is the Maximum Likelihood estimate θ of θ? 2. Show that the maximum likelihood estimator of θ is unbiased. 3. Were looking to cstimate the variance θ (1-9) of Xi . x being the empirical average 2(1-2). Check that T is not unli ator propose an unbiased estimator of θ(1-0).
Answer #1

1.

The likelihood equation is,

L( heta) = prod_{i=1}^{n}f(x_i; heta) = prod_{i=1}^{n} heta^{x_i}(1- heta)^{1-x_i} = heta^{sum_{i=1}^{n}x_i}(1- heta)^{n-sum_{i=1}^{n}x_i}

The log-likelihood function is,

logL( heta) = log heta sum_{i=1}^{n}x_i + log(1- heta) left ( n-sum_{i=1}^{n}x_i ight )

The maximum likelihood estimate is,

rac{partial }{partial heta}logL( heta) = sum_{i=1}^{n}x_i / heta - left ( n-sum_{i=1}^{n}x_i ight ) /(1- heta) = 0

Rightarrow rac{1- heta}{ heta} = left ( n /sum_{i=1}^{n}x_i ight ) - 1

rl 2-1

Rightarrow heta= left ( sum_{i=1}^{n}x_i /n ight ) = ar{x}

Thus, maximum likelihood estimate of heta is,

hat{ heta}= left ( sum_{i=1}^{n}x_i /n ight ) = ar{x}

2.

Eleft [ hat{ heta} ight ]= Eleft ( sum_{i=1}^{n}x_i /n ight ) =E[ar{x}] = heta (By central limit theorem and mean of Bernoulli distribution is heta)

Since, Eleft [ hat{ heta} ight ]= heta ,   hat{ heta} is unbiased estimate of heta.

3.

E[T] = E[ar{x}(1-ar{x})] = E[ar{x} - ar{x}^2] = E[ar{x}] - E[ar{x}^2]

= E[ar{x}] - (Var(ar{x})+E[ar{x}]^2) = heta - left ( heta(1- heta)/n + heta^2 ight )

= heta - left ( ( heta- heta^2)/n + heta^2 ight ) = (n heta- heta+ heta^2 - n heta^2)/n

= ((n-1) heta- (n-1) heta^2 )/n

= (n-1) heta(1 - heta)/n

Since, E[T] e heta(1 - heta) T is not unbiased estimator of heta(1 - heta) .

Let T' = (n/n-1) ar{x}(1 - ar{x})

Then,

T' = (n/n-1) ar{x}(1 - ar{x}) = (n/n-1) T

E[T'] = E[(n/n-1) T] = (n/n-1)E[ T]

= (n/n-1) (n-1) heta(1 - heta)/n = heta(1 - heta)

Thus,  T' is an unbiased estimator of heta(1 - heta) .